# Quasi-perfect blowups detect regularity

### v1

@@ Abstract.
This work gives a new characterization of regular Noetherian schemes (more generally algebraic spaces) as those for which blowups at closed points are quasi-perfect. In particular, regularity is controlled by the property that alterations or modifications of the scheme (or algebraic space) are quasi-perfect. Additionally, we study the local behavior of quasi-perfect morphisms, showing that for proper morphisms this property can be detected via stalks, and completions and (strict) Henselizations of those. Furthermore, we show that the locus of points where a proper morphism is quasi-perfect forms a Zariski open set.
@@ 1. Introduction
The zoo of singularities consists of many exotic creatures. To this day the study of its inhabitants is an active area of research. A very useful and helpful way to understand them is through geometrical and homological techniques. Typically, one associates to each inhabitant invariants. Such an invariant could be a category; it is natural to then try and study the information contained within, its deeper structure.
An important category associated to any Noetherian scheme $X$ is its derived category $D_{\operatorname{qc}}(X)$ of complexes of $\mathcal{O}_{X}$-modules with quasi-coherent cohomology. This ‘big’ category contains two important (essentially) small subcategories. Firstly, the category of perfect complexes, denoted $\operatorname{Perf}(X)$, and secondly the category of bounded complexes with coherent cohomology, denoted $D^{b}_{\operatorname{coh}}(X)$. One has the containment $\operatorname{Perf}(X)\subseteq D^{b}_{\operatorname{coh}}(X)$ which controls the regularity of the scheme: these categories coincide if and only if the scheme is regular. Furthermore, it is the latter which contains the skyscraper sheaves at closed points and it is the inclusion of these in the former, i.e. the skyscraper sheaves being perfect, which controls regularity at the corresponding points. It is this fact that leads to the main results in this work.
Probing $D^{b}_{\operatorname{coh}}(X)$ using categorical incarnations of geometric operations allows one to study its structure more precisely and uncover more information contained within. A key example is utilizing the derived pushforward along a proper morphisms $\mathbf{R}f_{\ast}\colon D^{b}_{\operatorname{coh}}(Y)\to D^{b}_{\operatorname{coh}}(X)$ to characterize certain singularities. Important examples include
  • rational singularities [Art66, Lip69] and their ‘mild’ variants [Lan25],
  • Du Bois singularities [DB81, Ste81, Ste83],
  • derived splinters [Bha12] and their ‘birational’ variants (following Kovács) [Lyu22, Lan25].
Such singularities are often considered to be ‘tame’ as they usually enjoy favorable properties; e.g. derived splinters being normal or Cohen–Macaulay in many settings. Loosely speaking, these type of singularities represent ‘slight deviations’ away from regularity in regards to (cohomological) behavior.
The story of this note began with [MS20]. There the authors showed that a local domain $R$, which is essentially of finite type over a field, is regular if and only if for every alteration $f\colon X\to\operatorname{Spec}(R)$ from a regular scheme one has $\mathbf{R}f_{\ast}\mathcal{O}_{X}\in\operatorname{Perf}(X)$. An analogue of this result in mixed characteristic was posed as a question (see [MS20, Question 1.1]). Our initial goal was to answer Ma and Schwede’s question. Whilst we were unsuccessful in that avenue, in the process it became clear that one can characterize regularity by having ‘enough probing morphisms’ be quasi-perfect in the sense of Lipman–Neeman [LN07]. Here quasi-perfect means that the derived pushforward preserves perfect complexes (see Definition 2.2).
As this characterization has, to the best of our knowledge, not appeared in the literature before, it felt worthwhile to write this short note to bring it to peoples attention. In fact, all of the results below are in the generality of algebraic spaces; but, we restrict to schemes in this introduction for ease.
Theorem A (see Theorem 3.3 ).
@@
A Noetherian scheme $X$ is regular if, and only if, the blowup of $X$ along any closed point is a quasi-perfect morphism.
Although A does not answer [MS20, Question 1.1], it provides new and interesting characterizations of regularity: one can characterize regularity by probing with alterations or modifications.
Corollary B (see Corollary 3.6).
An integral Noetherian scheme $X$ is regular if, and only if, every alteration (resp. modification) of $X$ is quasi-perfect.
Notably, we do not require the prior existence of an alteration or modification from a regular scheme. Therefore, it offers a ‘resolution free’ characterization of regularity; however, see Remark 3.7 for how one could show regularity using the existence of a ‘quasi-perfect resolution’. As a further application of A, we give a new proof that regularity descends along fpqc covers (see Corollary 3.11).
In order to prove A we use generation techniques for triangulated categories. Such techniques have been very successful in (cohomologically) characterizing singularities; for instance, the types of singularities mentioned above [LV25, LMV25, Lan25]. The crucial observation, in our case, is that one can use the skyscraper sheaf of a closed point to detect regularity at that point (we also extend this to algebraic spaces in Lemma 3.2, which seems to not have been written down before).
As A and B show that one can use quasi-perfectness of proper morphism to detect regularity, a natural next question is whether one can ‘practically’ detect quasi-perfectness. To this end, we show that this can be checked stalk-wise on the target, that is a proper morphisms is quasi-perfect if and only if all the base changes to the local rings are quasi-perfect. In particular, checking quasi-perfectness of proper morphisms reduces to showing this for morphisms with target the spectrum of a local ring $(R,\mathfrak{m})$. Furthermore, we show that one can replace $R$ by either its $\mathfrak{m}$-adic completion $\widehat{R}$ or the (strict) Henselization $R^{h}$ (resp. $R^{sh}$); neither affects the quasi-perfectness. Thus, all in all we have the following, our last main result.
Theorem C (see Theorem 3.14).
@@
Let $f\colon Y\to X$ be a proper morphism of schemes with $X$ separated and Noetherian. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#},\widehat{\mathcal{O}}_{X,\#},\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
@@
    the base change of $f$ to $\operatorname{Spec}(A_{p})$ is quasi-perfect for all $p\in X$.
In order to prove this theorem, we prove some lemmas, of perhaps independent interest, that allow one to check quasi-perfectness using specific types of flat covers (see Lemmas 3.9 and 3.10). Furthermore, as a corollary we show that the locus of points $p\in X$ for which the base change to $\mathcal{O}_{X,p}$ is quasi-perfect (or equivalently base change to the completion or (strict) Hensalization of the local ring) forms a Zariski open subset of $X$, see Corollary 3.15. An interesting application (which to the best of our knowledge was previously unknown) is that the property of the regular locus containing a nonempty open subset, i.e. $X$ being JJ-0, descends along proper surjective morphisms quite generally (we have a mild reducedness condition), see Proposition 3.17 for the details.
Let us also say a word concerning the restriction to a proper morphism in C. In view of A this is mild as the morphisms appearing there are proper. In fact, for a separated morphisms of finite type between Noetherian schemes, it being quasi-perfect implies that it is proper, see e.g. [Nee21, Proposition 0.20]; so in practice it is very mild indeed. However, the real reason we need to restrict to this is that in the proofs we need to leverage certain local-to-global techniques which require working with bounded and coherent objects (e.g. [AIL10, Theorem 4.1]).
We end this introduction with a few words on the generalization to algebraic spaces. Although we stated the above results for schemes, these are proven below for algebraic spaces. This is done Section 3 after a short recollection of necessary definitions and notation in Section 2. We expect the more general results to be beneficial for future work for those studying singularities of algebraic spaces which is recently been gaining traction. For example, recent progress has been made in establishing the minimal model program for excellent algebraic spaces in characteristic zero, see [LM22].
@@ 2.1. Generation
Let $\mathcal{T}$ be a triangulated category. A triangulated subcategory of $\mathcal{T}$ is called thick if it is closed under direct summands (and so automatically strict, i.e. closed under isomorphisms). For any collection of objects $S$ of $\mathcal{T}$, let $\operatorname{add}(S)$ denote the smallest strictly full subcategory of $\mathcal{T}$ containing $S$ that is closed under shifts, finite coproducts and direct summands. Inductively define
@@
An important feature of this construction is that $\langle S\rangle:=\cup_{n\geq 0}\langle S\rangle_{n}$ is the smallest thick triangulated subcategory of $\mathcal{T}$ containing $S$. Observe also that it is possible to rewrite the definition of $\langle\mathcal{S}\rangle_{n}$ using $\operatorname{Hom}(\langle\mathcal{S}\rangle_{1},\langle\mathcal{S}\rangle_{n-1})$ instead (so there is no choice ‘left vs. right’).
@@ 2.2. Algebraic Spaces
Our conventions follow that of the Stacks Project [Sta25]. Throughout we work with algebraic spaces in the ‘absolute’ setting, i.e. over the base scheme $\operatorname{Spec}(\mathbb{Z})$; but, of course, everything remains valid when working over an arbitrary base.
Let $X$ be an algebraic space throughout this subsection . We denote by $|X|$ the set of points of $X$, which obtains a topology by giving it the quotient topology via an étale surjective morphism from a scheme; see e.g. [Sta25, Tag 03BT]. The algebraic spaces that are considered in this work will always be decent in the sense of [Sta25, Tag 03I7]. This condition ensures that there is a good relationship between the topology of $|X|$ and properties of $X$ itself. Note that quasi-separated algebraic spaces are automatically decent; in fact, for locally Noetherian algebraic spaces being descent is equivalent to being quasi-separated [Sta25, Tag 0BB6].
Defining irreducibility for algebraic spaces is a bit subtle as this notion is not étale local. Therefore, recall that an algebraic space $X$ is called irreducible when its underlying topological spaces $|X|$ is irreducible; it is not clear whether one can characterize this using coverings from schemes (e.g. using étale surjective morphisms from schemes). An integral algebraic spaces, then, is one that is irreducible, reduced and decent [Sta25, Tag 0AD3].
@@
Let $X_{\acute{e}t}$ denote the small étale site of $X$. The assignment $U\mapsto\Gamma(U,\mathcal{O}_{U})$ defines a sheaf of rings on $X_{\acute{e}t}$, denoted $\mathcal{O}_{X}$ and called the structure sheaf (of $X$). Consequently, it makes sense to speak of sheaves $\mathcal{O}_{X}$-modules on $X_{\acute{e}t}$; denote the (Grothendieck) Abelian category of such sheaves by $\operatorname{Mod}(X)$. The Abelian subcategories of quasi-coherent resp. coherent $\mathcal{O}_{X}$-modules are denoted by $\operatorname{Qcoh}(X)$ resp. $\operatorname{Coh}(X)$.
@@
Any morphisms $f\colon Y\to X$ of algebraic spaces induces a morphism of ringed topoi [Sta25, Tag 03G8], which similarly as in the case of schemes leads to a derived adjunction $\mathbf{L}f^{\ast}\colon D(X)\rightleftarrows D(Y)\colon\mathbf{R}f_{\ast}$; see e.g. [Sta25, Tag 07A6]. When $f$ is quasi-compact and quasi-separated, the derived pushforward preserves objects with quasi-coherent cohomology (see e.g. [Sta25, Tag 08FA]) and so the adjunction restricts to $D_{\operatorname{qc}}$, i.e. we have an adjunction $\mathbf{L}f^{\ast}\colon D_{\operatorname{qc}}(X)\rightleftarrows D_{\operatorname{qc}}(Y)\colon\mathbf{R}f_{\ast}$.
@@
Perfect complexes can be defined for any ringed site [Sta25, Tag 08G4], and so, in particular, for $X$ by looking at its small étale site. Recall that a complex is strictly perfect if it is a bounded complex with each term a direct summand of a finite free module. A complex is perfect if it is locally strictly perfect. The subcategory of $D_{\operatorname{qc}}(X)$ consisting of perfect complexes is denoted by $\operatorname{Perf}(X)$; this is a thick subcategory.
When $X$ is quasi-compact and quasi-separated, the perfect complexes are exactly the compact objects of $D_{\operatorname{qc}}(X)$, see [Sta25, Tag 09M8]. Recall that an object in a triangulated category with arbitrary direct sums is said to be compact if the covariant Hom out of that object commutes with direct sums. Moreover, in this case, $D_{\operatorname{qc}}(X)$ is compactly generated by a single perfect complex, see [Sta25, Tag 0AEC]. Here $P$ compactly generating means that it is compact and any non-zero object admits a non-zero morphism from $P$, any such object is called a compact generator. It is worthwhile noting that $P$ being a compact generator for $D_{\operatorname{qc}}(X)$ is equivalent to it being a classical generator for $\operatorname{Perf}(X)$, i.e.  $\operatorname{Perf}(X)=\langle P\rangle$. This follows as $D_{\operatorname{qc}}(X)$ is compactly generated, see e.g. [Sta25, Tag 09SR].
Lastly, for future reference, let us note the following lemma which shows that one can check perfectness flat locally; see e.g. [HR17, Lemma 4.1] or [DLMRP25, Lemma 2.2]. For the definition of the the fpqc topology, see [Sta25, Tag 03MP]; important examples of fpqc covers are fppf and so also smooth and étale covers.
@@
Let $\overline{x}$ be a geometric point of $X$, i.e. $\overline{x}$ is a morphism from the spectrum of an algebraically closed field. For any presheaf $\mathcal{F}$ on $X_{\acute{e}t}$ one defines the stalk of $\mathcal{F}$ at $\overline{x}$ as
$\mathcal{F}_{\bar{x}}:=\mathop{\mathrm{colim}}\nolimits_{(U,\overline{u})}\mathcal{F}(U)$
where the colimit runs over the étale neighborhoods of $\overline{x}$ in $X$. Taking stalks at geometric points are exactly the points of the small étale site of $X$ (see e.g. [Sta25, Tag 04K6]). The étale local ring of $X$ at $\overline{x}$ is by definition the stalk of the structure sheaf at $\overline{x}$, i.e. $\mathcal{O}_{X,\overline{x}}:=(\mathcal{O}_{X})_{\overline{x}}$. This is in fact a local ring (as can be seen from the following paragraph).
To make the latter more explicit, pick an étale neighborhood $(U,\overline{u})$ of $\overline{x}$ where $U$ a scheme. Then, from [Sta25, Tag 04KF], we have isomorphisms
@@
where $\mathcal{O}_{U,u}^{sh}$ is the strict Henselization of $\mathcal{O}_{U,u}$ and $u\in U$ is the point at which $\overline{u}$ is centered. We refer to [Sta25, Tag 0BSK] for the definition of (strict) Henselizations of local rings. Because of Equation 2.1, for arbitrary $x\in|X|$, by abuse of notation we sometimes denote $\mathcal{O}_{X,x}^{sh}:=\mathcal{O}_{X,\overline{x}}$ where $\overline{x}$ is any geometric point lying over $x$.
@@
where now the colimit is taken over the ‘elementary étale neighborhoods’ (see [Sta25, Tag 0BGU] for the definition). As in the previous paragraph one can make this more explicit in terms of a covering. Indeed, by [Sta25, Tag 0EMY], picking an elementary etale neighborhood $(U,u)\to(X,x)$ one has $\mathcal{O}_{X,x}^{h}\cong\mathcal{O}_{U,u}^{h}$, where $\mathcal{O}_{U,u}^{h}$ is the Henselization of the local ring $\mathcal{O}_{U,u}$ of $U$ at $u$. In particular,
@@
Pick a point $x\in|X|$. When $X$ is decent there exists, by [Sta25, Tag 03K4], a unique monomorphism $\operatorname{Spec}(k)\to X$ representing $x$. The field $k$ appearing is called the residue field of $X$ at $x$ and is denoted $\kappa(x)$.
@@
Recall that $X$ is regular if there exists an étale surjective morphism from a regular scheme, see e.g. [Sta25, Tag 03E5]. Furthermore, a point $x\in|X|$ is regular if there exists an étale morphism $f\colon U\to X$ from a scheme and a point $u\in U$ such that $f(u)=x$ and the stalk $\mathcal{O}_{U,u}$ is a regular local ring; this is well-defined by [Sta25, Tag 0AH7]. It is straightforward to see that an algebraic space is regular if and only if all of its points are regular. Moreover, by [Sta25, Tag 0AHA], a point $x\in|X|$ is regular if and only if the étale local ring $\mathcal{O}_{X,\overline{x}}$ is a regular local ring for any (equivalently some) geometric point $\overline{x}$ lying over $x$.
When $X$ is a Noetherian algebraic space, it follows from [DLMRP25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$. Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLMRP25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e. $\operatorname{Perf}(X)=\langle G\rangle_{n+1}$ for some $G\in\operatorname{Perf}(X)$ and $n\geq 0$.

### v2

# A note on quasi-perfect morphisms

@@ Abstract.
This note records two results concerning quasi-perfect morphisms between Noetherian algebraic spaces. First, we give a new characterization of regular Noetherian algebraic spaces as those for which blowups at closed points are quasi-perfect. Secondly, we study the local behavior of quasi-perfect proper morphisms. In particular, we show that quasi-perfectness of a proper morphism can be detected at their étale local rings, completions, and (strict) Henselizations. As a corollary, the locus of points where a proper morphism is quasi-perfect is Zariski open.
@@ 1. Introduction
Given a quasi-compact quasi-separated morphism $f\colon Y\to X$ of schemes, Neeman’s Brown Representability Theorem provides a right adjoint $f^{\times}$ to the derived pushforward $\mathbf{R}f_{\ast}$ on $D_{\operatorname{qc}}$. While $f^{\times}$ preserves small coproducts for complexes that are uniformly homologically bounded below, this property fails in general. Motivated by [VER69, Lemma 1 & Corollary 2], Lipman–Neeman [LN07] introduced the notion of quasi-perfect morphisms for quasi-compact quasi-separated schemes to remedy this. The defining property is that $f^{\times}$ preserves small coproducts on all of $D_{\operatorname{qc}}$ or, equivalently, that $\mathbf{R}f_{\ast}$ preserve perfect complexes; see [LN07, Theorem 1.2 & Proposition 2.1]. Whilst there are numerous examples of quasi-perfect morphisms—for instance, any proper morphisms of finite Tor dimension between Noetherian schemes—quasi-perfectness generally fails. For example, this is the case for closed immersions into a non-regular Noetherian scheme.
More recent developments have recast these ideas in a categorical framework. For example, Balmer–Dell’Ambrogio–Sanders develop a duality theory for rigidly-compactly generated tensor-triangulated categories [BDS16, Theorem 1.7]. However, when interpreted in the setting of schemes, this perspective is largely global in nature. In particular, the local behavior of quasi-perfectness remains less understood; for instance, how quasi-perfectness at the local rings of the target influences quasi-perfectness globally.
In this short note we make two novel observations regarding quasi-perfect morphisms between algebraic spaces. The first observation is how one can use these morphisms to characterize regularity.
Proposition (Proposition˜3.3).
@@
A Noetherian algebraic space $X$ is regular if and only if the blowup of $X$ along any closed point is a quasi-perfect morphism.
To the best of our knowledge, this characterization has not appeared before in the literature even for schemes. The simple, but crucial, observation is that a closed point on a Noetherian algebraic space is regular if and only if the blowup at that point is quasi-perfect. This follows from the fact that perfectness of the skyscraper sheaf at a closed point detects regularity of that point; we prove this in Lemma˜3.1, which seems to not have been written down for algebraic spaces before and might be of independent interest. Furthermore, as an application of Proposition˜3.3, we give a new proof that regularity descends along fpqc covers (see Corollary˜4.5).
The second observation concerns the local behavior of quasi-perfectness for proper morphisms. This is summarized in the following result.
Theorem (Theorem˜4.8).
@@
Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
@@
Theorem˜4.8 provides a mechanism for detecting quasi-perfectness by reducing the question to targets that are local rings. The proof relies on several auxiliary lemmas, potentially of independent interest, which allow quasi-perfectness to be checked using specific classes of flat covers (see Lemmas˜4.2 and 4.3). As a corollary, we show that the locus of points $p$, for which the base change to one of the classes of local rings $A_{p}$ from Theorem˜4.8 is quasi-perfect, forms a Zariski open subset.
Corollary (Corollary˜4.9).
Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜4.8. Then the locus
$S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
is an open subset of $X$ or $U$ depending on the choice of $A_{\#}$.
One can think of $S$ as the ‘the quasi-perfect locus of $f$’. The fact it is always open for proper morphisms appears to be new even for schemes. As an application, we show that the property that the regular locus contains a nonempty open subset (i.e. that $X$ is JJ-0) descends along proper surjective morphisms under mild hypotheses, which are satisfied if, for example, the target is reduced (see Proposition˜4.11).
Finally, we comment on the restriction to proper morphisms in Theorem˜4.8. In view of Proposition˜3.3, this is mild since the morphisms appearing there are proper. In fact, for a separated morphism of finite type between Noetherian schemes, quasi-perfectness already implies properness (see [NEE21, Proposition 0.20]); so this is quite a mild restriction. The properness assumption is required in our arguments to employ local-to-global techniques involving bounded coherent objects (e.g. [AIL10, Theorem 4.1]).
We expect the more general results to be useful in future work on singularities of algebraic spaces, a topic that has recently been gaining traction. For example, recently the minimal model program has been established for excellent algebraic spaces in characteristic zero [LM22].
@@ 2.1. Generation
Let $\mathcal{T}$ be a triangulated category. Recall that a strictly full subcategory is a full subcategory closed under isomorphisms. A triangulated subcategory of $\mathcal{T}$ is called thick if it is closed under direct summands (and so automatically strict, i.e. closed under isomorphisms). For any collection of objects $S$ of $\mathcal{T}$, let $\operatorname{add}(S)$ denote the smallest strictly full subcategory of $\mathcal{T}$ containing $S$ that is closed under shifts, finite coproducts and direct summands. Following [BV03], inductively, define
@@
An important feature of this construction is that $\langle S\rangle:=\cup_{n\geq 0}\langle S\rangle_{n}$ is the smallest thick triangulated subcategory of $\mathcal{T}$ containing $S$. Observe also that it is possible to rewrite the definition of $\langle\mathcal{S}\rangle_{n}$ using $\operatorname{Hom}(\langle\mathcal{S}\rangle_{1},\langle\mathcal{S}\rangle_{n-1})$, i.e. there is no choice ‘left vs. right’.
@@ 2.2. Algebraic Spaces
Our conventions follow that of the Stacks Project [STA26]. Throughout we work with algebraic spaces in the ‘absolute’ setting, i.e. over the base scheme $\operatorname{Spec}(\mathbb{Z})$. Of course, everything remains valid over an arbitrary base.
Let $X$ be an algebraic space throughout this subsection. We denote by $|X|$ the set of points of $X$, endowed with the quotient topology via any étale surjective morphism from a scheme; see e.g. [STA26, Tag 03BT]. The algebraic spaces that are considered in this work will always be decent in the sense of [STA26, Tag 03I7]. This condition ensures that there is a good relationship between the topology of $|X|$ and properties of $X$ itself. Note that quasi-separated algebraic spaces are automatically decent; in fact, for locally Noetherian algebraic spaces being descent is equivalent to being quasi-separated [STA26, Tag 0BB6].
Defining irreducibility for algebraic spaces is a bit subtle as this notion is not étale local. Therefore, recall that an algebraic space $X$ is called irreducible when its underlying topological spaces $|X|$ is irreducible; it is not clear whether one can characterize this using coverings from schemes (e.g. using étale surjective morphisms from schemes). An algebraic spaces is integral if it’s irreducible, reduced, and decent [STA26, Tag 0AD3].
@@
Let $X_{\acute{e}t}$ denote the small étale site of $X$. The assignment $U\mapsto\Gamma(U,\mathcal{O}_{U})$ defines a sheaf of rings on $X_{\acute{e}t}$, denoted $\mathcal{O}_{X}$ and called the structure sheaf (of $X$). Consequently, it makes sense to speak of sheaves $\mathcal{O}_{X}$-modules on $X_{\acute{e}t}$; denote the (Grothendieck) Abelian category of such sheaves by $\operatorname{Mod}(X)$. The Abelian subcategories of quasi-coherent (resp. coherent) $\mathcal{O}_{X}$-modules are denoted by $\operatorname{Qcoh}(X)$ (resp. $\operatorname{Coh}(X)$).
@@
Any morphisms $f\colon Y\to X$ of algebraic spaces induces a morphism of ringed topoi [STA26, Tag 03G8], which leads to a derived adjunction $\mathbf{L}f^{\ast}\colon D(X)\rightleftarrows D(Y)\colon\mathbf{R}f_{\ast}$ as in the case of schemes; see e.g. [STA26, Tag 07A6]. When $f$ is quasi-compact and quasi-separated, the derived pushforward preserves objects with quasi-coherent cohomology (see e.g. [STA26, Tag 08FA]) and so the adjunction restricts to $D_{\operatorname{qc}}$, i.e. we have an adjunction $\mathbf{L}f^{\ast}\colon D_{\operatorname{qc}}(X)\rightleftarrows D_{\operatorname{qc}}(Y)\colon\mathbf{R}f_{\ast}$.
@@
Perfect complexes can be defined for any ringed site [STA26, Tag 08G4], and so, in particular, for $X$ by looking at its small étale site. Recall that a complex is strictly perfect if it is a bounded complex with each term a direct summand of a finite free module. A complex is perfect if it is locally strictly perfect. The subcategory of $D_{\operatorname{qc}}(X)$ consisting of perfect complexes is denoted by $\operatorname{Perf}(X)$; this is a thick subcategory.
When $X$ is quasi-compact and quasi-separated, the perfect complexes are exactly the compact objects of $D_{\operatorname{qc}}(X)$, see [STA26, Tag 09M8]. Recall that an object in a triangulated category with arbitrary direct sums is said to be compact if the covariant Hom out of that object commutes with direct sums. Moreover, in this case, $D_{\operatorname{qc}}(X)$ is compactly generated by a single perfect complex, see [STA26, Tag 0AEC]. Here, an object $P$ compactly generating means that it is compact and any non-zero object admits a non-zero morphism from some shift of $P$. Any such object is called a compact generator. It is worthwhile noting that $P$ being a compact generator for $D_{\operatorname{qc}}(X)$ is equivalent to it being a classical generator for $\operatorname{Perf}(X)$, i.e. $\operatorname{Perf}(X)=\langle P\rangle$. This follows as $D_{\operatorname{qc}}(X)$ is compactly generated, see e.g. [STA26, Tag 09SR].
Lastly, for future reference, let us note the following lemma which shows that one can check perfectness flat locally; see e.g. [HR17, Lemma 4.1] or [DLR+25, Lemma 2.2]. For the definition of the the fpqc topology, see [STA26, Tag 03MP]; important examples of fpqc covers are fppf, and hence, also smooth and étale covers.
@@
Let $\overline{x}$ be a geometric point of $X$, i.e. $\overline{x}$ is a morphism from the spectrum of an algebraically closed field to $X$. For any presheaf $\mathcal{F}$ on $X_{\acute{e}t}$, one defines the stalk of $\mathcal{F}$ at $\overline{x}$ as $\mathcal{F}_{\bar{x}}:=\mathop{\mathrm{colim}}\nolimits_{(U,\overline{u})}\mathcal{F}(U)$ where the colimit runs over the étale neighborhoods of $\overline{x}$ in $X$. The stalks at geometric points defined this way are exactly the stalks at points of the small étale site of $X$ (see e.g. [STA26, Tag 04K6]). The étale local ring of $X$ at $\overline{x}$ is by definition the stalk of the structure sheaf at $\overline{x}$, i.e. $\mathcal{O}_{X,\overline{x}}:=(\mathcal{O}_{X})_{\overline{x}}$. This is in fact a local ring (as can be seen from the following paragraph).
To make the latter more explicit, pick an étale neighborhood $(U,\overline{u})$ of $\overline{x}$ where $U$ a scheme. Then, from [STA26, Tag 04KF], we have isomorphisms
@@
where $\mathcal{O}_{U,u}^{sh}$ is the strict Henselization of $\mathcal{O}_{U,u}$ and $u\in U$ is the point at which $\overline{u}$ is centered. We refer to [STA26, Tag 0BSK] for the definition of (strict) Henselizations of local rings. Because of (2.1), for arbitrary $x\in|X|$, by abuse of notation we sometimes denote $\mathcal{O}_{X,x}^{sh}:=\mathcal{O}_{X,\overline{x}}$ where $\overline{x}$ is any geometric point lying over $x$.
@@
where now the colimit is taken over the ‘elementary étale neighborhoods’ (see [STA26, Tag 0BGU] for the definition). As in the previous paragraph, one can make this more explicit in terms of a covering. Indeed, by [STA26, Tag 0EMY], picking an elementary etale neighborhood $(U,u)\to(X,x)$ one has $\mathcal{O}_{X,x}^{h}\cong\mathcal{O}_{U,u}^{h}$, where $\mathcal{O}_{U,u}^{h}$ is the Henselization of the local ring $\mathcal{O}_{U,u}$ of $U$ at $u$. In particular,
@@
Pick a point $x\in|X|$. When $X$ is decent there exists, by [STA26, Tag 03K4], a unique monomorphism $\operatorname{Spec}(k)\to X$ representing $x$. The field $k$ appearing is called the residue field of $X$ at $x$ and is denoted $\kappa(x)$.
@@
Recall that $X$ is regular if there exists an étale surjective morphism from a regular scheme, see e.g. [STA26, Tag 03E5]. Furthermore, a point $x\in|X|$ is regular if there exists an étale morphism $f\colon U\to X$ from a scheme and a point $u\in U$ such that $f(u)=x$ and the stalk $\mathcal{O}_{U,u}$ is a regular local ring; this is well-defined by [STA26, Tag 0AH7]. It is straightforward to see that an algebraic space is regular if and only if all of its points are regular. Moreover, by [STA26, Tag 0AHA], a point $x\in|X|$ is regular if and only if the étale local ring $\mathcal{O}_{X,\overline{x}}$ is a regular local ring for any (equivalently some) geometric point $\overline{x}$ lying over $x$.
When $X$ is a Noetherian algebraic space, it follows from [DLR+25, Theorem 3.7] that regularity of $X$ is equivalent to $D^{b}_{\operatorname{coh}}(X)=\operatorname{Perf}(X)$. Furthermore, when $X$ is a separated Noetherian algebraic space of finite Krull dimension, by [DLR+25, Theorem 5.1], regularity of $X$ is equivalent to $\operatorname{Perf}(X)$ being strongly generated, i.e. $\operatorname{Perf}(X)=\langle G\rangle_{n+1}$ for some $G\in\operatorname{Perf}(X)$ and $n\geq 0$.